### v2

# Mean-field approach to finite-size fluctuations in the Kuramoto-Sakaguchi model

Abstract
We develop an ab initio approach to describe the statistical behavior of finite-size fluctuations in the deterministic Kuramoto-Sakaguchi model. We obtain explicit expressions for the covariance function of fluctuations of the complex order parameter and determine the variance of its magnitude entirely in terms of the equation parameters. Our results rely on an explicit complex-valued formula for the solution of the Adler equation. We present analytical results for both the sub- and the super-critical case. Moreover, our framework does not require any prior knowledge about the structure of the partially synchronized state. We corroborate our results with numerical simulations of the full Kuramoto-Sakaguchi model. The proposed methodology is sufficiently general such that it can be applied to other interacting particle systems.
phase oscillators, Kuramoto-Sakaguchi model, finite-size fluctuations, mean-field approach
Introduction. Synchronization is a generic phenomenon inherent in interacting oscillators that determines their ability to exhibit a wide range of complex dynamical behavior [27, 48, 41, 1, 2, 13, 44]. It has been observed in a plethora of natural and engineered systems, including pace-maker cells of circadian rhythms [53], networks of neurons [3], chemical oscillators [25, 50] and power grids [15]. The main features of synchronization are captured by the paradigmatic Kuramoto-Sakaguchi (KS) model [45]
$\dot{\psi}_{i}=\omega_{i}+\frac{\displaystyle K}{\displaystyle N}\sum\limits_{j=1}^{N}\sin(\psi_{j}-\psi_{i}-\lambda)$ (1)
for the collective behavior of a system of $N$ all-to-all coupled phase oscillators. In this model, $\omega_{i}$ are natural frequencies drawn randomly and independently from a given distribution $g(\omega)$ and $K$ and $\lambda$ are real parameters quantifying the coupling strength and the phase lag, respectively. It is often convenient to rewrite Eq. (1) as
$\dot{\psi_{i}}=\omega_{i}+K\>\mathrm{Im}\left(Z(t)e^{-i\psi_{i}}e^{-i\lambda}\right),$
which reveals that each oscillator interacts with the other oscillators only via the mean field
$Z(t)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i\psi_{j}(t)},$ (2)
which is also coined the complex order parameter and quantifies the degree of synchronization.
A fully synchronized state with $\theta_{1}=\theta_{2}=\dots=\theta_{N}$ is characterized by $r(t)=|Z(t)|=1$, whereas a completely disordered (incoherent) state is characterized by $r(t)=\mathcal{O}(1/\sqrt{N})$. It is well known that for sufficiently large values of the coupling strength $K$ the system becomes fully synchronized, whereas for sufficiently small values of $K$ the oscillators behave independently and form an incoherent state. For intermediate values of the coupling strength, partially synchronized states are observed in which $r(t)$ fluctuates around a constant value between 0 and $1$. On a microscopic level the oscillator population splits into two groups: coherent oscillators $\mathcal{C}$, which are phase locked, and rogue oscillators $\mathcal{R}$, which drift with respect to the mean field.
In the thermodynamic limit of infinitely many oscillators each oscillator interacts with the other oscillators via a constant collective mean. This limit is well understood and described by mean-field theory [45, 33, 34, 35]. In particular, the rotation frequency of the coherent oscillators $\Omega_{\infty}$ and the constant order parameter $r_{\infty}$ are determined by a self-consistency relation, first derived in [45] and later formulated in a more convenient complex form in [33],
$\frac{\displaystyle 1}{\displaystyle K}e^{i\lambda}=i\int_{-\infty}^{\infty}g(\Omega_{\infty}+Kr_{\infty}s)h(s)ds,$ (3)
where
$h(s)=\left\{\begin{array}[]{lcl}(1-\sqrt{1-s^{-2}})s&\mbox{ for }&|s|>1\\[2.84526pt] s-i\sqrt{1-s^{2}}&\mbox{ for }&|s|\leq 1.\end{array}\right$. (4)
However, recently finite-size effects have gained much attention. Numerical experiments have led to a better understanding of finite-size effects and provided clear evidence for their significance [9, 10, 7, 32, 22, 47, 55, 49]. In particular, several finite-size phenomena were observed such as a stochastic drift of oscillators which disappears in the thermodynamic limit [16, 30, 18, 19] and emergent random chimera switching in a deterministic two-population KS model [24]. Capturing finite-size effects analytically is a notoriously hard problem. Relaxation rates and pair correlation functions were obtained using kinetic theory [20, 5, 54] and the finite-size scaling near criticality has been successfully quantified [23, 22, 21, 37]. Recently, substantial progress has been made for stochastic systems of coupled oscillators [52, 4]. However, analytical methods for deterministic systems still remain underdeveloped. A general scheme for analyzing fluctuations in deterministic systems of coupled phase oscillators was proposed in [26, 11, 8]. In particular, accurate approximations were obtained in the subcritical case of the KS model (1) with $\lambda=0$ and a Lorentzian frequency distribution [8]. However, this approach relies on several approximations which do not carry over to the case of more general frequency distributions such as Gaussians as we will show (see Appendix K). In this Letter we will capture finite-size effects in a first principle way and provide explicit expressions for the covariance function and the variance of the order parameter.
Our theory is based on the assumption of an underlying central limit theorem describing fluctuations around the mean-field theory limit. This implies that fluctuations in the deterministic KS model decay with $1/\sqrt{N}$ and constitute a Gaussian process determined entirely by its mean and covariance function (see Appendix F for numerical evidence for this assumption). This points to the range of validity of our theory: we require sufficiently large system size $N$ in order to allow for the validity of the central limit theorem.
Specifically, we consider the modified order parameter
$Z_{\mathrm{mod}}(t)=Z(t)\frac{\overline{Z_{\mathrm{c}}(t)}}{|Z_{\mathrm{c}}(t)|}\quad\mbox{with}\quad Z_{\mathrm{c}}(t)=\frac{1}{N}\sum\limits_{j\in\mathcal{C}}e^{i\psi_{j}(t)},$
where the bar denotes the complex-conjugate. Compared to the original order parameter (2), the modified order parameter is represented in the frame of reference moving with the synchronized cluster. This ensures that, contrary to $Z(t)$, the modified order parameter does not exhibit phase diffusion [42, 40] (see Appendix I). Recently, it was numerically recognized that fluctuations
$\zeta(t)=\sqrt{N}(Z_{\mathrm{mod}}(t)-\langle Z_{\mathrm{mod}}(t)\rangle)$ (5)
around the thermodynamic mean $\langle Z_{\mathrm{mod}}(t)\rangle$, where $\langle\cdot\rangle$ denotes time averaging, can be approximated by a Gaussian process [55, 24]. Fluctuations in $Z_{\mathrm{mod}}(t)$, with a variance that decays as $1/N$ and vanishes in the thermodynamic limit, lead to fluctuations in other macroscopic variables, including the order parameter with fluctuations
$\delta(t)=\sqrt{N}(|Z_{\mathrm{mod}}(t)|-\langle|Z_{\mathrm{mod}}(t)|\rangle)=\sqrt{N}(r(t)-\langle r(t)\rangle).$ (6)
Gaussian processes are entirely determined by their mean, accessible via classical mean-field theory, and their covariance function
$R(\tau)=\langle\zeta(t)\overline{\zeta(t+\tau)}\rangle.$ (7)
Previous work [55, 24] has shown the efficacy of such an approach, however, these works relied on approximating the general Gaussian process by an Ornstein-Uhlenbeck process the parameters of which were estimated only numerically and needed to be recalculated for each set of equation parameters. Moreover, a precise a priori knowledge of which oscillators are synchronized and which are not was required.
In this Letter, we find explicit analytical expressions for the covariance function (7) and the variance of finite-size fluctuations of the order parameter $V=\langle\delta^{2}(t)\rangle$, given entirely in terms of equation parameters $K$, $\lambda$ and $g(\omega)$. We present expressions for both the partially synchronized and the completely incoherent state. Remarkably, our expressions do not require any a priori knowledge about which oscillators partake in the synchronized cluster and which do not. This opens up the way to perform stochastic model reduction for deterministic systems of coupled oscillators where the drift and diffusion coefficients are given entirely in terms of the microscopic equation parameters. Such reductions have so far been restricted to multi-scale systems [17].
Our main results can be formulated as follows. For partially synchronized states we predict
$R_{\mathrm{ps}}(\tau)=Kr_{\infty}\int\limits_{-\infty}^{\infty}\frac{\displaystyle(\overline{h^{2}(s)}-1)(1-|h^{2}(s)|)g(\Omega_{\infty}+Kr_{\infty}s)}{\displaystyle\overline{h^{2}(s)}-\exp(iKr_{\infty}s\sqrt{1-s^{-2}}\>\tau)}ds$ (8)
and
$V_{\mathrm{ps}}=\frac{\displaystyle 1}{\displaystyle 2}R_{\mathrm{ps}}(0)+\mathcal{O}\left(\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\right)\quad\mbox{for}\quad N\gg 1.$ (9)
For completely incoherent states we find
$R_{\mathrm{incoh}}(\tau)=\int_{-\infty}^{\infty}g(\omega)e^{-i\omega\tau}d\omega$ (10)
and
$\sqrt{N}\langle r(t)\rangle=\frac{\displaystyle\sqrt{\pi}}{\displaystyle 2}\quad\mbox{and}\quad V_{\mathrm{incoh}}=1-\frac{\displaystyle\pi}{\displaystyle 4}\quad\mbox{for}\quad N\gg 1.$ (11)
Our analytical results and their efficacy in capturing finite-size fluctuations are illustrated in Figs. 1 and 2, where we compare our predictions with simulations of the full KS model (1) for a network of $1,000$ oscillators with a Gaussian natural frequency distribution $g(\omega)$ with unit-variance. Fig. 1 shows the covariance function $R(\tau)$ for a partially synchronized state described by (8) and for a completely incoherent state described by (10). Fig. 2(a) shows the averaged order parameter $\langle r(t)\rangle$ as a function of the coupling strength $K$. The order parameter of the full KS model (1) is very well reproduced by classical mean-field theory (3) for values of the coupling strength $K\geq K_{\rm{crit}}$. Moreover, our extension of the mean-field results to the completely incoherent state given by (11) reproduces the order parameter for the subcritical range $K<K_{\rm{crit}}$. Fig. 2(b) shows that the variance of the order parameter is well captured by our analytical expressions (9) and (11) for coupling strength away from the bifurcation at $K_{\rm{crit}}$. Additional numerical results and the description of our numerical setup are provided in Appendix H and Appendix E, respectively.
Figure 1: Covariance $R(\tau)$ of fluctuations $\zeta(t)$ of the complex order parameter for the KS model (1). (Top row) Partially synchronized state at $K=6$ and $\lambda=\pi/4$, with theoretical prediction (8). (Bottom row) Completely incoherent state at $K=0.5$ and $\lambda=\pi/4$, with theoretical prediction (10).
Figure 2: (a) Averaged order parameter $\langle r(t)\rangle$, with theoretical predictions from classical mean-field theory (3) for $K>K_{\rm{crit}}\approx 1.92$ and from our approach (11) for $K<K_{\rm{crit}}$, and (b) its variance $V$, with theoretical predictions (9), (11), versus coupling strength $K$ in the KS model (1) with $\lambda=\pi/4$.
Covariance function for a partially synchronized state. Our approach is based on the following prerequisites. (i) Moving into the corotating frame $\theta_{i}(t)=\psi_{i}(t)-\arg Z_{\mathrm{c}}(t)$ and neglecting the assumed Gaussian fluctuations for $N\to\infty$, Eq. (1) becomes the so called Adler equation
$\dot{\theta}_{i}=\omega_{i}-\Omega_{\infty}+K\>\mathrm{Im}\left(Z_{\infty}e^{-i\theta_{i}}e^{-i\lambda}\right).$ (12)
(ii) We use the well-known fact (see Appendix A for details) that the probability density of the solution of the Adler equation (12) is given by the Ott-Antonson ansatz [35]
$f_{\phi}(\theta,\omega)=\frac{\displaystyle g(\omega)}{\displaystyle 2\pi}\left\{1+\sum\limits_{n=1}^{\infty}\left[\overline{z}_{\phi}^{n}(\omega)e^{in\theta}+z_{\phi}^{n}(\omega)e^{-in\theta}\right]\right\},$ (13)
where
$z_{\phi}(\omega)=e^{i\phi}h\left(\frac{\displaystyle\omega-\Omega_{\infty}}{\displaystyle Kr_{\infty}}\right).$ (14)
It follows directly from (13) that for every non-negative integer $n$ we have
$\int_{0}^{2\pi}\frac{\displaystyle f_{\phi}(\theta,\omega)}{\displaystyle g(\omega)}e^{in\theta}d\theta=z_{\phi}^{n}(\omega),$ (15)
which, upon employing (2) and (3), implies
$Z_{\infty}=\int_{-\infty}^{\infty}d\omega\int_{0}^{2\pi}f_{\phi}(\theta,\omega)e^{i\theta}d\theta=-r_{\infty}ie^{i\lambda}e^{i\phi}.$ (16)
Moreover, if $\theta_{i}(t)$ is a solution of the scalar Adler equation (12), then for any $2\pi$-periodic function $F(\theta)$ we have the ergodicity property
$\langle F(\theta_{j}(t))\rangle=\int_{0}^{2\pi}\frac{\displaystyle f_{\phi}(\theta,\omega_{j})}{\displaystyle g(\omega_{j})}F(\theta)d\theta.$ (17)
(iii) Finally, we establish an explicit solution of the Adler equation (12). Introducing $\Delta_{j}=\omega_{j}-\Omega_{\infty}$ and $H_{\infty}=KZ_{\infty}e^{-i\lambda}$, solutions of the Adler equation $\theta_{j}(t)$ with initial conditions $\theta_{j}(0)$ are given for $|\Delta_{j}|\neq|H_{\infty}|$ by
$e^{i\theta_{j}(t)}=W(t,H_{\infty},\Delta_{j},e^{i\theta_{j}(0)}),$ (18)
with
$W(t,H,\Delta,\sigma)=Ha_{1}(t)+\frac{\displaystyle a_{2}(t)\sigma}{\displaystyle 1+\overline{H}a_{1}(t)\sigma},$ (19)
where
$a_{1}(t)=i(e^{i\xi t}-1)/\chi_{1}(t)\quad\mbox{and}\quad a_{2}(t)=4\xi^{2}e^{i\xi t}/\chi_{1}^{2}(t)$
with $\chi_{1}(t)=(\Delta-\xi)e^{i\xi t}-(\Delta+\xi)$ and
$\xi=\left\{\begin{array}[]{lcl}-i\sqrt{|H|^{2}-\Delta^{2}}&\mbox{ for }&|\Delta|\leq|H|,\\[5.69054pt] \hphantom{-i}\sqrt{\Delta^{2}-|H|^{2}}&\mbox{ for }&|\Delta|>|H|.\end{array}\right$. (20)
Note that for $|\Delta_{j}|<|H_{\infty}|$, the $j$th oscillator is coherent and the dynamics of $\theta_{j}(t)$ converges to a fixed point as $t\to\infty$, while for $|\Delta_{j}|>|H_{\infty}|$, the oscillator is incoherent with underlying periodic dynamics. Details on the derivation of (18), (19) are provided in Appendix B.
The covariance function (7) can be expressed in terms of the fluctuations (5) as
$R_{\mathrm{ps}}(\tau)$ $=$ $N\langle Z_{\mathrm{mod}}(t)\overline{Z_{\mathrm{mod}}(t+\tau)}\rangle-N|Z_{\infty}|^{2}$ (21)
$=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j,k=1}^{N}\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle-N|Z_{\infty}|^{2}.$
If either $j\in\mathcal{C}$ or $k\in\mathcal{C}$ the phase asymptotically approaches a fixed point and the summands can be written as
$\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle=\langle e^{i\theta_{j}(t)}\rangle\langle e^{-i\theta_{k}(t)}\rangle.$ (22)
Moreover, the independence identity (22) also holds if $j,k\in\mathcal{R}$ and $\sqrt{\Delta_{j}^{2}-|H_{\infty}|^{2}}$ and $\sqrt{\Delta_{k}^{2}-|H_{\infty}|^{2}}$ are incommensurable (see Appendix J), and hence generically the sum in (21) can be written as
$\sum\limits_{j,k=1}^{N}\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle=\sum\limits_{j,k=1}^{N}\langle e^{i\theta_{j}(t)}\rangle\langle e^{-i\theta_{k}(t)}\rangle$
$+\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{-i\theta_{j}(t+\tau)}\rangle-\sum\limits_{j\in\mathcal{R}}\left|\langle e^{i\theta_{j}(t)}\rangle\right|^{2}.$
Substitution into (21) yields, upon using $\langle Z_{\mathrm{mod}}(t)\rangle=Z_{\infty}$,
$R_{\mathrm{ps}}(\tau)=\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{-i\theta_{j}(t+\tau)}\rangle-\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\left|\langle e^{i\theta_{j}(t)}\rangle\right|^{2}.$ (23)
Now, using (17) to replace temporal averages by averages over the phases and the explicit formula (18) for $e^{i\theta_{j}(t+\tau)}$, we obtain
$\langle e^{i\theta_{j}(t)}\rangle$ $=$ $z_{\phi}(\omega_{j}),$
$\left\langle e^{i\theta_{j}(t)}e^{-i\theta_{j}(t+\tau)}\right\rangle$ $=$ $\int_{0}^{2\pi}e^{i\theta}\overline{W}(\tau,H_{\infty},\Delta_{j},e^{i\theta})\times$
$\times\frac{\displaystyle f_{\phi}(\theta,\omega_{j})}{\displaystyle g(\omega_{j})}d\theta$ $=$ $\frac{\displaystyle\overline{Ha_{1}(\tau)}}{\displaystyle\overline{z_{\phi}(\omega_{j})}}(|z_{\phi}(\omega_{j})|^{2}-1)+1.$
(Note that for the second correlation, we used identity (15) after expanding the solution (18)–(19) in a geometric series. Details of the calculations can be found in Appendix B.) Altogether, this yields
$R_{\mathrm{ps}}(\tau)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\left[\frac{\displaystyle\overline{Ha_{1}(\tau)}}{\displaystyle\overline{z_{\phi}(\omega_{j})}}-1\right]\left(|z_{\phi}(\omega_{j})|^{2}-1\right).$
Finally, in the thermodynamic limit, averaging over indices $j$ can be replaced by averaging over the distribution $g(\omega)$, so we can write
$R_{\mathrm{ps}}(\tau)=\int\limits_{-\infty}^{\infty}\left[\frac{\displaystyle\overline{Ha_{1}(\tau)}}{\displaystyle\overline{z_{\phi}(\omega)}}-1\right]\left(|z_{\phi}(\omega)|^{2}-1\right)g(\omega)d\omega.$ (24)
After some algebraic manipulations (see Appendix C), the covariance function (24) simplifies to our expression (8). Note that if $|\omega-\Omega_{\infty}|\leq|H_{\infty}|=Kr_{\infty}$, $|z_{\phi}|=1$. Therefore such a set of $\omega$ does not contribute to the integral $R_{\mathrm{ps}}(\tau)$. This has the desirable consequence that in order to evaluate the covariance function, no prior knowledge is required about which oscillators are and which are not synchronized.
Covariance function for a completely incoherent state. Our analysis of the completely incoherent state is based on the following prerequisites. (i) In this state, the phases $\psi_{j}$ tend to be uniformly distributed on the interval $0,2\pi]$, and their probability density in the thermodynamic limit is given by
$f_{\mathrm{incoh}}(\psi,\omega,t)=\frac{\displaystyle g(\omega)}{\displaystyle 2\pi}.$ (25)
(ii) The global order parameter shows no collective oscillations, and $Z(t)=0$ as $N\to\infty$. This implies that the dynamics of each oscillator is approximately described by $\dot{\psi}_{i}=\omega_{i}$, and therefore $\psi_{i}(t)=\omega_{i}t+\psi_{i}(0)$.
Consider the fluctuations
$\zeta_{\mathrm{incoh}}(t)=\sqrt{N}(Z(t)-\langle Z(t)\rangle)$
of the complex order parameter in this incoherent state. The corresponding covariance function is then given by
$R_{\mathrm{incoh}}(\tau)$ $=$ $\langle\zeta_{\mathrm{incoh}}(t)\overline{\zeta_{\mathrm{incoh}}(t+\tau)}\rangle$
$=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j,k=1}^{N}\left\langle e^{i\psi_{j}(t)}e^{-i\psi_{k}(t+\tau)}\right\rangle-N|\langle Z(t)\rangle|^{2}.$
For different incommensurable frequencies $\omega_{j}$ and $\omega_{k}$, the variables $e^{i\psi_{j}(t)}$ and $e^{i\psi_{k}(t)}$ are independent, which implies again an identity of the form (22), and we obtain
$R_{\mathrm{incoh}}(\tau)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\left\langle e^{i\psi_{j}(t)}e^{-i\psi_{j}(t+\tau)}\right\rangle-\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\left|\langle e^{i\psi_{j}(t)}\rangle\right|^{2}.$
Moreover, using the ergodicity property and the stationary density for incoherent oscillators (25), we find
$\langle e^{i\psi_{j}(t)}\rangle=\int_{0}^{2\pi}e^{i\psi}\frac{\displaystyle f_{\mathrm{incoh}}(\psi,\omega_{j},t)}{\displaystyle g(\omega_{j})}d\psi=0$ (26)
and
$\langle e^{i\psi_{j}(t)}e^{-i\psi_{j}(t+\tau)}\rangle=e^{-i\omega_{j}\tau}.$
Finally, in the large-$N$ limit we can replace averaging over indices $j$ by averaging over the distribution $g(\omega)$ and arrive at our expression (10) for the covariance function $R_{\rm{incoh}}(\tau)$. Note that due to the normalization condition for $g(\omega)$, we have $\langle\left|\zeta_{\mathrm{incoh}}(t)\right|^{2}\rangle=R_{\mathrm{incoh}}(0)=1$.
Variance of the order parameter for a partially synchronized state. Rewriting (5) as
$Z_{\mathrm{mod}}(t)=\langle Z_{\mathrm{mod}}(t)\rangle+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\zeta(t)=Z_{\infty}+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\zeta(t)$
implies
$|Z_{\mathrm{mod}}(t)|^{2}=|Z_{\infty}|^{2}+\frac{\displaystyle 2}{\displaystyle\sqrt{N}}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\frac{\displaystyle 1}{\displaystyle N}|\zeta(t)|^{2},$
which upon expansion for $N\gg 1$ becomes
$|Z_{\mathrm{mod}}(t)|=|Z_{\infty}|+\frac{\displaystyle 1}{\displaystyle\sqrt{N}|Z_{\infty}|}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\mathcal{O}\left(\frac{\displaystyle 1}{\displaystyle N}\right).$
This allows to express the fluctuations of the order parameter (6) as
$\delta(t)=\frac{\displaystyle 1}{\displaystyle|Z_{\infty}|}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\mathcal{O}\left(\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\right),$
and the variance of the order parameter becomes
$V_{\mathrm{ps}}=\frac{\displaystyle 1}{\displaystyle 2}\mathrm{Re}\>\left(\frac{\displaystyle\overline{Z_{\infty}^{2}}}{\displaystyle|Z_{\infty}|^{2}}\tilde{R}_{\mathrm{ps}}(0)\right)+\frac{\displaystyle 1}{\displaystyle 2}R_{\mathrm{ps}}(0)+\mathcal{O}\left(\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\right),$
where $\tilde{R}_{\rm{ps}}(t)$ denotes the pseudo-covariance $\tilde{R}_{\rm{ps}}(t)=\langle\zeta(t)\zeta(t+\tau)\rangle$. Employing calculations similar to those for $R_{\mathrm{ps}}(t)$ (see Appendix D for details), we can show that $\tilde{R}_{\mathrm{ps}}(\tau)=0$ for all $\tau\geq 0$. The vanishing of the pseudo-covariance can be intuitively understood by evoking a symmetry between the real and imaginary parts of the fluctuations $\zeta$ and their independence. Using $\tilde{R}_{\mathrm{ps}}(\tau)=0$ for all $\tau\geq 0$, leads to our result (9).
Variance of the order parameter for a completely incoherent state. In the large-$N$ limit, for a completely incoherent state we have $\langle Z(t)\rangle=0$, and therefore $|Z(t)|=|\zeta_{\mathrm{incoh}}(t)|/\sqrt{N}.$ Using the definition of $\zeta_{\mathrm{incoh}}(t)$, we write
$\zeta_{\mathrm{incoh}}(t)=\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\sum\limits_{j=1}^{N}\cos\psi_{j}(t)+\frac{\displaystyle i}{\displaystyle\sqrt{N}}\sum\limits_{j=1}^{N}\sin\psi_{j}(t).$
In the completely incoherent state all phases $\psi_{j}(t)$ are uniformly distributed in $0,2\pi]$. This implies that in the large-$N$ limit fluctuations satisfy a central limit theorem, and $\zeta_{\mathrm{incoh}}(t)$ is a two-dimensional random vector distributed according to a bivariate normal distribution. The mean and variances are readily determined. For each $t\geq 0$ the function $\cos\psi_{j}(t)$ is a random variable with zero expected value $\langle\cos\psi_{j}\rangle_{\mathrm{ens}}=0$ and variance
$\left\langle\cos^{2}\psi_{j}\right\rangle_{\mathrm{ens}}=\frac{\displaystyle 1}{\displaystyle 2\pi}\int_{0}^{2\pi}\cos^{2}\psi\>d\psi=\frac{\displaystyle 1}{\displaystyle 2},$
where $\langle\cdot\rangle_{\mathrm{ens}}$ denotes the ensemble average. Similarly, $\sin\psi_{j}(t)\sim\mathcal{N}(0,1/2)$. Moreover, the random variables $\cos\psi_{j}(t)$ and $\sin\psi_{j}(t)$ are uncorrelated since
$\langle\cos\psi_{j}\sin\psi_{j}\rangle_{\mathrm{ens}}=\frac{\displaystyle 1}{\displaystyle 2\pi}\int_{0}^{2\pi}\cos\psi\sin\psi\>d\psi=0.$
Hence $\mathrm{Re}\>(\zeta_{\mathrm{incoh}}(t))\sim\mathcal{N}(0,1/2)$ and $\mathrm{Im}\>(\zeta_{\mathrm{incoh}}(t))\sim\mathcal{N}(0,1/2)$, and $\left|\zeta_{\mathrm{incoh}}(t)\right|$ has a Rayleigh distribution with expected value
$\langle\left|\zeta_{\mathrm{incoh}}(t)\right|\rangle=\sqrt{\pi}/2$
and variance
$V_{\mathrm{incoh}}=\langle\left|\zeta_{\mathrm{incoh}}(t)\right|^{2}\rangle-\langle\left|\zeta_{\mathrm{incoh}}(t)\right|\rangle^{2}=1-\pi/4.$
Note that the value $\langle\left|\zeta_{\mathrm{incoh}}(t)\right|^{2}\rangle=1$, which follows from the last two identities, agrees with the value $\langle\left|\zeta_{\mathrm{incoh}}(t)\right|^{2}\rangle=R_{\mathrm{incoh}}(0)$ obtained above.
Discussion and outlook. In this Letter, we have proposed a fully analytical approach to capture the statistical behavior of finite-size fluctuations in the deterministic KS model (1). In particular, our approach provides
(i) expressions for the covariance function of fluctuations of the complex order parameter and the variance of the order parameter entirely in terms of the KS model parameters,
(ii) expressions for both sub- and super-critical coupling strengths,
(iii) expressions that do not require prior knowledge about which oscillators are synchronized and which are not.
Our theory relies on the assumption of an underlying central limit theorem governing the Gaussian statistics of finite-size fluctuations. Remarkably, the statistical independence of fluctuations of distinct oscillators, which were essential for finding explicit expressions, follows directly from our solution of the Adler equation.
Simulations for the full KS model showed that our theory performs well away from criticality. On the other hand, our approximations were less accurate when approaching the critical coupling strength $K_{\rm{crit}}$, demarcating the transition from a desynchronized to a partially synchronized state. The explicit expressions for the covariance function rely on the explicit solution of the Adler equation. Hence, the quality of our theory crucially depends on the Adler equation being a good approximation of the full system. This is only guaranteed if fluctuations can be neglected in Eq. (1), and breaks down when approaching criticality (see Appendix G for details). We envisage extensions of our framework to the near critical case using a perturbative approach around the expressions derived in this work.
Our approach is sufficiently general to allow for further applications to other interacting particle systems such as $\theta$-neuron models and more complex network topologies going beyond the all-to-all coupling considered in this Letter. The mathematical tools developed here can potentially be adapted to study finite-size fluctuations in noisy coupled oscillator systems [31, 29, 19, 52, 43, 12, 56, 4] finite-size sampling effects [39, 38, 14] and dynamics of complexified KS models [51, 28]. Further, it is known that the incoherent state and partially synchronized states in the KS model are weakly chaotic attractors, where the chaoticity can be attributed to the presence of finite-size fluctuations. Therefore, it is natural to expect that knowing the characteristics of finite-size fluctuations, may aid in computing finite-size corrections of Lyapunov exponents [6].
In summary, our results fully specify the statistical behavior of finite-size fluctuations in the KS model which have been numerically shown to be Gaussian processes. This opens up the way to perform systematic stochastic model reductions for such interacting particle systems from first principles and to construct reduced stochastic equations for designated collective variables [17, 46].
Acknowledgements
The work of O.E.O. was partly supported by the Deutsche Forschungsgemeinschaft under Grant No. OM 99/2-3. O.E.O acknowledges the financial support and hospitality of the Sydney Mathematical Research Institute (SMRI), where parts of this work was completed on an International Visitor Program. The authors are grateful for insightful discussions with A. Pikovsky.
Appendix A Thermodynamic limit for the KS model (1)
Most of the dynamical states found in the KS model (1) with finite $N$, have complex chaotic behavior. However, in the thermodynamic limit $N\to\infty$, they admit a simple analytic representation. If $N\gg 1$, the state of the phase oscillators $\{\psi_{i}(t)\}$ can be described by a probability density function $f(\psi,\omega,t)$, which obeys the continuity equation
$\frac{\displaystyle\partial f}{\displaystyle\partial t}+\frac{\displaystyle\partial}{\displaystyle\partial\psi}(fv)=0,$ (27)
where
$v(\psi,\omega,t)=\omega+\frac{\displaystyle K}{\displaystyle 2i}\left[e^{-i\lambda}\mathcal{Z}(t)e^{-i\psi}-e^{i\lambda}\overline{\mathcal{Z}}(t)e^{i\psi}\right]$
is the continuum version of the velocity field in Eq. (1), and
$\mathcal{Z}(t)=\int_{-\infty}^{\infty}d\omega\int_{0}^{2\pi}f(\psi,\omega,t)e^{i\psi}d\psi$ (28)
is the continuum version of the complex order parameter (2).
It is well-known [35, 36] that the long-term dynamics of the continuity equation (27) asymptotically approaches the Ott-Antonsen manifold for $t\to\infty$, consisting of distributions of the form
$f(\psi,\omega,t)=\frac{\displaystyle g(\omega)}{\displaystyle 2\pi}\left\{1+\sum\limits_{n=1}^{\infty}\left[\overline{z}^{n}(\omega,t)e^{in\psi}+z^{n}(\omega,t)e^{-in\psi}\right]\right\},$
where $z(\omega,t)$ satisfies the inequality $|z|\leq 1$ and solves the integro-differential equation
$\frac{\displaystyle\partial z}{\displaystyle\partial t}=i\omega z(\omega,t)+\frac{\displaystyle K}{\displaystyle 2}e^{-i\lambda}\mathcal{G}z-\frac{\displaystyle K}{\displaystyle 2}e^{i\lambda}z^{2}(\omega,t)\mathcal{G}\overline{z}$ (29)
with the integral operator
$(\mathcal{G}z)(t)=\int_{-\infty}^{\infty}g(\omega)z(\omega,t)d\omega.$
Remarkably, all statistically stationary states of the KS model (1) lie on the Ott-Antonsen manifold [33, 34]. In particular, the zero solution $z(\omega,t)=0$ of Eq. (29) corresponds to the completely incoherent state in (1), while all stationary partially synchronized states in (1) are represented by rotating waves
$z(\omega,t)=h\left(\frac{\displaystyle\omega-\Omega_{\infty}}{\displaystyle Kr_{\infty}}\right)e^{i\Omega_{\infty}t},$ (30)
where $h(s)$ is defined by (4) and $(r_{\infty},\Omega_{\infty})\in(0,1)\times\mathbb{R}$ is a pair of numbers satisfying the self-consistency equation (3).
The physical meaning of the parameters $r_{\infty}$ and $\Omega_{\infty}$ can be understood by inserting the probability density function $f(\psi,\omega,t)$ corresponding to (30) into Eq. (28). This yields
$\mathcal{Z}(t)=-r_{\infty}ie^{i\lambda}e^{i\Omega_{\infty}t},$ (31)
and therefore
$|\mathcal{Z}(t)|=r_{\infty}$ (32)
is the magnitude of the order parameter, while $\Omega_{\infty}$ is the angular speed of its phase.
If in the KS model (1) we move into the corotating frame
$\theta_{i}(t)=\psi_{i}(t)-\Omega_{\infty}t+\phi\quad\mbox{with some}\quad\phi\in\mathbb{R},$ (33)
the probability density function in the new frame will take on the time-independent form (13), which is used for the calculations in the main text. Importantly, for an appropriately chosen value of $\phi$, we can guarantee that (16) yields the thermodynamic limit of the modified order parameter $Z_{\mathrm{mod}}(t)$.
Appendix B Explicit solutions of the Adler equation
Let us consider the Adler equation written in complex form,
$\dot{\theta}=\Delta+\mathrm{Im}\>\left(He^{-i\theta}\right)=\Delta+\frac{\displaystyle He^{-i\theta}-\overline{H}e^{i\theta}}{\displaystyle 2i}$ (34)
with arbitrary $\Delta\in\mathbb{R}$ and $H\in\mathbb{C}$.
Proposition 1
If $|\Delta|\neq|H|$, the solution of the Adler equation (34) is of the form
$e^{i\theta(t)}=W(t,H,\Delta,e^{i\theta(0)}):=\frac{\displaystyle iH(e^{i\xi t}-1)-\chi_{2}(t)e^{i\theta(0)}}{\displaystyle i\overline{H}(e^{i\xi t}-1)e^{i\theta(0)}+\chi_{1}(t)},$
where $\xi$ is defined in Eq. (20), and
$\chi_{1}(t)=(\Delta-\xi)e^{i\xi t}-(\Delta+\xi),\quad\chi_{2}(t)=(\Delta+\xi)e^{i\xi t}-(\Delta-\xi).$
In addition, in the degenerate case $|\Delta|=|H|$, the solution of the Adler equation (34) is given by
$e^{i\theta(t)}=W_{0}(t,H,\Delta,e^{i\theta(0)}):=z_{*}+\frac{\displaystyle e^{i\theta(0)}-z_{*}}{\displaystyle 1+(e^{i\theta(0)}-z_{*})\overline{H}t/2},$
where $z_{*}=i\Delta/H$. (For brevity of notation, we do not specify that $\xi$, $\chi_{1}(t)$, $\chi_{2}(t)$ and $z_{*}$ depend on $\Delta$ and $H$.)
Proof: Multiplying Eq. (34) by $ie^{i\theta}$, we obtain a complex differential equation for $z=e^{i\theta}$:
$\dot{z}=\frac{\displaystyle 1}{\displaystyle 2}H+i\Delta z-\frac{\displaystyle 1}{\displaystyle 2}\overline{H}z^{2}.$ (35)
For $|\Delta|\neq|H|$, the right-hand side of this equation can be written as
$\frac{\displaystyle 1}{\displaystyle 2}H+i\Delta z-\frac{\displaystyle 1}{\displaystyle 2}\overline{H}z^{2}=-\frac{\displaystyle 1}{\displaystyle 2}\overline{H}(z-z_{+})(z-z_{-}),$
where $z_{\pm}=i(\Delta\pm\xi)/\overline{H}$. Using the method of separation of variables and the initial condition $z(0)=\sigma$, we obtain
$z(t)=\frac{\displaystyle z_{+}(\sigma-z_{-})e^{i\xi t}-z_{-}(\sigma-z_{+})}{\displaystyle(\sigma-z_{-})e^{i\xi t}-(\sigma-z_{+})},$
or equivalently $e^{i\theta(t)}=W(t,H,\Delta,e^{i\theta(0)})$.
In the degenerate case $|\Delta|=|H|$, we have $z_{+}=z_{-}=z_{*}$. Integrating Eq. (35), we obtain $e^{i\theta(t)}=W_{0}(t,H,\Delta,e^{i\theta(0)})$.
Remark 1
In the case $|\Delta|\neq|H|$, we can write
$W(t,H,\Delta,\sigma)=Ha_{1}(t)+\frac{\displaystyle a_{2}(t)\sigma}{\displaystyle 1+\overline{H}a_{1}(t)\sigma},$
with
$a_{1}(t)=\frac{\displaystyle i(e^{i\xi t}-1)}{\displaystyle\chi_{1}(t)}\quad\mbox{and}\quad a_{2}(t)=\frac{\displaystyle 4\xi^{2}e^{i\xi t}}{\displaystyle\chi_{1}^{2}(t)}.$
Moreover, in this case, we have $|Ha_{1}(t)|<1$, and hence
$W(t,H,\Delta,\sigma)=Ha_{1}(t)+a_{2}(t)\sum\limits_{n=0}^{\infty}(-1)^{n}\overline{H}^{n}a_{1}^{n}(t)\sigma^{n+1}$
is an absolutely convergent series for $|\sigma|\leq 1$.
Remark 2
If $(r_{\infty},\Omega_{\infty})\in(0,1)\times\mathbb{R}$ satisfy the self-consistency equation (3), it can be easily verified that $z_{\varphi}(\omega)$ defined by (14) is a fixed point of the Adler equation (34) with $\Delta=\omega-\Omega_{\infty}$ and $H=-iKr_{\infty}e^{i\phi}$. Therefore,
$W\left(t,-iKr_{\infty}e^{i\phi},\omega-\Omega_{\infty},z_{\phi}(\omega)\right)=z_{\phi}(\omega)$
for all $t,\omega,\phi\in\mathbb{R}$.
Remark 3
Under the assumptions of Remark 2, using Remark 1 and formula (15), we can show
[…]
$\left.+\overline{a_{2}(t)}\sum\limits_{n=0}^{\infty}(-1)^{n}H^{n}\overline{a_{1}^{n}(t)}e^{-in\theta}\right]\frac{\displaystyle f_{\phi}(\theta,\omega)}{\displaystyle g(\omega)}d\theta$
$=\overline{Ha_{1}(t)}z_{\phi}(\omega)+\overline{a_{2}(t)}\sum\limits_{n=0}^{\infty}(-1)^{n}H^{n}\overline{a_{1}^{n}(t)}\overline{z}_{\phi}^{n}(\omega)$
$=\overline{Ha_{1}(t)}\left(z_{\phi}(\omega)-\frac{\displaystyle 1}{\displaystyle\overline{z_{\phi}(\omega)}}\right)+\frac{\displaystyle 1}{\displaystyle\overline{z_{\phi}(\omega)}}\overline{W(t,H,\Delta,z_{\phi}(\omega))}$
$=\frac{\displaystyle\overline{Ha_{1}(t)}}{\displaystyle\overline{z_{\phi}(\omega)}}(|z_{\phi}(\omega)|^{2}-1)+1.$
Appendix C Derivation of Eq. (8) from Eq. (24)
Using (4) it is easy to verify that
$|H|\,h\left(\frac{\displaystyle\Delta}{\displaystyle|H|}\right)=\left\{\begin{array}[]{lcl}\Delta-\xi\>\mathrm{sgn}(\Delta)&\mbox{ for }&|\Delta|>|H|,\\[5.69054pt] \Delta+\xi&\mbox{ for }&|\Delta|\leq|H|,\end{array}\right$. (36)
where $\xi$ is given by (20). Note that formula (36) holds for all $\Delta\in\mathbb{R}$ and $H\in\mathbb{C}$.
Now, we consider the integrand of Eq. (24), recalling that $H=KZ_{\infty}e^{-i\lambda}=-iKr_{\infty}e^{i\phi}$. Recognizing that in this case
$\frac{\displaystyle H}{\displaystyle z_{\phi}(\omega)}=-\frac{\displaystyle iKr_{\infty}e^{i\phi}}{\displaystyle e^{i\phi}h\left(\frac{\displaystyle\omega-\Omega_{\infty}}{\displaystyle Kr_{\infty}}\right)}=-\frac{\displaystyle i|H|}{\displaystyle h\left(\frac{\displaystyle\Delta}{\displaystyle|H|}\right)}$
does not depend on $\phi$, we find
$\frac{\displaystyle Ha_{1}(\tau)}{\displaystyle z_{\phi}(\omega)}-1=-\frac{\displaystyle i|H|}{\displaystyle h\left(\frac{\displaystyle\Delta}{\displaystyle|H|}\right)}\frac{\displaystyle i(e^{i\xi t}-1)}{\displaystyle\chi_{1}(t)}-1$
$=\frac{\displaystyle|H|^{2}}{\displaystyle|H|h\left(\frac{\displaystyle\Delta}{\displaystyle|H|}\right)}\frac{\displaystyle e^{i\xi t}-1}{\displaystyle(\Delta-\xi)e^{i\xi t}-(\Delta+\xi)}-1.$ (37)
For $\Delta>|H|$, upon using (36) and the identity $\Delta^{2}-\xi^{2}=|H|^{2}$, (37) is written as
$\frac{\displaystyle Ha_{1}(\tau)}{\displaystyle z_{\phi}(\omega)}-1=\frac{\displaystyle|H|^{2}(e^{i\xi t}-1)}{\displaystyle(\Delta-\xi)^{2}e^{i\xi t}-(\Delta+\xi)(\Delta-\xi)}-1$
$=\frac{\displaystyle e^{i\xi t}-1}{\displaystyle h^{2}(\Delta/|H|)e^{i\xi t}-1}-1=\frac{\displaystyle 1-h^{2}(\Delta/|H|)}{\displaystyle h^{2}(\Delta/|H|)-e^{-i\xi t}}.$
Similarly, for $\Delta\leq|H|$ we obtain
$\frac{\displaystyle Ha_{1}(\tau)}{\displaystyle z_{\phi}(\omega)}-1=\frac{\displaystyle 1-h^{2}(\Delta/|H|)}{\displaystyle h^{2}(\Delta/|H|)-e^{i\xi t}},$
and thus in both cases we have
$\frac{\displaystyle Ha_{1}(\tau)}{\displaystyle z_{\phi}(\omega)}-1=\frac{\displaystyle 1-h^{2}(\Delta/|H|)}{\displaystyle h^{2}(\Delta/|H|)-e^{-i\Delta\sqrt{1-|H|^{2}/\Delta^{2}}t}}.$
Substituting into Eq. (24), we obtain Eq. (8).
[…]
Figure 3: Empirical histograms of the real (a) and imaginary (b) part of the fluctuations $\zeta$ and of $d[{\rm{arg}}Z_{\mathrm{c}}(t)]/dt$ centered around $\Omega_{\infty}\approx-1.71$ (c) for the KS model (1) with $N=1,000$ oscillators, phase lag $\lambda=\pi/4$, and $K=3$. The continuous lines (online red) show a best-fit Gaussian with the same variance.
[…]
Figure 4: Variance of the order parameter $V$ (a) and of the real (b) and imaginary (c) part of the fluctuations $\zeta$ as a function of the non-entrained oscillators $N_{r}$ for the KS model (1) with phase lag $\lambda=\pi/4$ and $K=3$ for $N=60,125,250,500,1000$. The continuous lines (online red) are lines of best fit indicating a scaling law of $N_{r}^{-0.96}$, $N_{r}^{-0.96}$ and $N_{r}^{-0.95}$, respectively.
Appendix D Pseudo-covariance function for a partially synchronized state
Using the definition of the pseudo-covariance function $\tilde{R}_{\rm{ps}}(t)=\langle\zeta(t)\zeta(t+\tau)\rangle$ and employing the established independence of the functions $e^{i\theta_{j}(t)}$ and $e^{i\theta_{k}(t)}$ for $j\neq k$, by analogy with (23), we obtain
$\tilde{R}_{\mathrm{ps}}(\tau)=\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{i\theta_{j}(t+\tau)}\rangle-\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}\rangle^{2}.$
Using the ergodicity property (17) and performing calculations as in Remark 3, we find
$\langle e^{i\theta_{j}(t)}e^{i\theta_{j}(t+\tau)}\rangle=\int\limits_{0}^{2\pi}e^{i\theta}W(\tau,H_{\infty},\Delta_{j},e^{i\theta})\frac{\displaystyle f_{\phi}(\theta,\omega_{j})}{\displaystyle g(\omega_{j})}d\theta$
$=z_{\phi}(\omega_{j})W(\tau,H_{\infty},\Delta_{j},z_{\phi}(\omega_{j}))=z_{\phi}^{2}(\omega_{j}).$
Recalling that $\langle e^{i\theta_{j}(t)}\rangle=z_{\phi}(\omega_{j})$, we obtain $\tilde{R}_{\mathrm{ps}}(\tau)=0$. Note that this relation is valid regardless of the choice of $\phi$ in the definition of the corotating frame (33).
Appendix E Details of the numerical setup
Given a distribution $g(\omega)$, we generate a set of $N$ natural frequencies $\omega_{j}$ according to
$\int_{-\infty}^{\omega_{j}}g(\omega)d\omega=\frac{\displaystyle j}{\displaystyle N+1},\qquad j=1,\dots,N.$
This equiprobable sampling avoids finite-size effects such as randomly clustered natural frequencies leading to small synchronized clusters with non-zero mean frequencies and their respective interactions [39, 14]. For each pair of parameters $K$ and $\lambda$, we perform simulations using a standard Runge-Kutta solver with constant time step $dt=0.02$. We start from initial conditions chosen randomly from the interval $0,2\pi]$, discard a transient period of $T_{\mathrm{transient}}=10^{5}$ time units and use the subsequent interval of length $T_{\mathrm{max}}=10^{5}$ time units to calculate statistical quantities such as means, variances, and covariances.
Appendix F Numerical evidence for the assumption on the Gaussianity of fluctuations
Previous work [55, 24] has shown numerically that finite-size fluctuations in the Kuramoto-Sakaguchi model (1) constitute a Gaussian process, consistent with an underlying central limit theorem. For completeness, we provide here further numerical evidence. Fig. 3(a) and (b) show the empirical histogram of the real and the imaginary part of the fluctuations $\zeta$ obtained from a long time simulation over $50,000$ time units with a sampling time of $dt=0.05$ of the KS model together with a best fit Gaussian clearly establishing the Gaussian character of the fluctuations. For $K=3$ we find a value of the skewness for the real and imaginary part of the fluctuations $\zeta$ of 0.058-0.058 and 0.051-0.051, respectively, and a kurtosis of $2.98$ for both. We further show in Fig. 3(c) the empirical histogram for the temporal derivative of the mean phase of the entrained oscillators $\tfrac{d}{dt}{\rm{arg}}Z_{c}(t)$, which also shows a clear Gaussian statistics, consistent with the phase diffusion discussed below.
Figure 5: Normalized covariance $R_{\rm{ps}}(\tau)/R_{\rm{ps}}(0)$ (real part only) of the KS model (1) for partially synchronized states with $N=1,000$ oscillators at $\lambda=\pi/4$ for various values of the coupling strength $K$.
To further establish numerical evidence for an underlying central limit theorem we show in Fig. 4 that the variance of the fluctuations $\zeta/\sqrt{N}$ scales with the number of non-entrained oscillators $N_{r}$ as $1/N_{r}$, as expected for Gaussian fluctuations described by an underlying central limit theorem. We show the scaling for the variance of the order parameter in Fig. 4(a) and of the real and imaginary part of the fluctuations $\zeta$ in Fig. 4 (b) and (c), respectively.
We remark that the central limit theorem holds for all values of the coupling strength $K>K_{\rm{crit}}$. For $K=1.98$ we find a value of the skewness for the real and imaginary part of the fluctuations $\zeta$ of 0.024-0.024 and 0.007-0.007, and a kurtosis of $2.98$ and $3.04$, respectively, suggestive of a Gaussian distribution. The Gaussian character of the empirical probability density function of the fluctuations alone, however, is not sufficient for the existence of a functional central limit theorem which is concerned with convergence to a stochastic process rather than to just random variables. We show in Figure 5 plots of the covariance function of the real part of the fluctuations, $\mathrm{Re}\,R_{\mathrm{ps}}(\tau)$ which decays to zero for all $K>K_{\rm{crit}}$ with the decay of correlation time increasing when approaching criticality.
Appendix G Limitations of the theory
From the previous Section it is clear that in order for the central limit theorem to be valid we require sufficiently large values of $N$. Indeed, for small values of $N=50$ we only have $N_{r}=6$ for $K=4$. Figure 6(a) and (b) clearly show that the empirical histograms of the real and imaginary parts of fluctuations $\zeta$ are far from Gaussian. The speed of the phase of the synchronized order parameter $d[{\rm{arg}}Z_{\mathrm{c}}(t)]/dt$ is still Gaussian, illustrating the phase diffusion discussed below.
[…]
Figure 6: Empirical histograms of the real (a) and imaginary (b) part of the fluctuations $\zeta$ and of $d[{\rm{arg}}Z_{\mathrm{c}}(t)]/dt$ centered around $\Omega_{\infty}\approx-2.39$ (c) for the KS model (1) with $N=50$ oscillators, phase lag $\lambda=\pi/4$, and $K=4$. The continuous lines (online red) show a best-fit Gaussian with the same variance.
Figure 7: Ratio of the standard deviation of the fluctuations $\sqrt{V}$ of the order parameter $r(t)$ and its mean $\langle r(t)\rangle$ for $N=1,000$ oscillators at $\lambda=\pi/4$ as a function of the coupling strength $K$.
The explicit expressions for the covariance functions crucially depends on the explicit solution of the Adler equation (12). To obtain the Adler equation we neglected fluctuations of the modified order parameter, replacing $Z_{\mathrm{mod}}(t)$ by its constant thermodynamic limit. This is only valid provided the fluctuations are small in size compared to the thermodynamic mean values. Approaching criticality, however, the size of fluctuations increases. Figure 7 shows the ratio of the standard deviation of the fluctuations $\sqrt{V}$ of the order parameter $r(t)$ and its mean as a function of the coupling strength $K$. It is clearly seen how fluctuations increase rapidly in size when approaching criticality at $K_{\rm{crit}}\approx 1.92$. This implies that the Adler equation and its solution become less valid closer to criticality. Hence, the quality of our theory depends on the value of the coupling strength $K$. Indeed, Figure 2(c) in the main text illustrates that our theory works well for coupling strengths sufficiently far from criticality, and gets worse the closer we approach criticality.
Appendix H Additional numerical results
To demonstrate how the accuracy of our method changes as we approach the critical coupling strength $K_{\mathrm{crit}}\approx 1.92$, we show in Figs. 811 additional results for partially synchronized states at $K=5,4,3,2$ for $\lambda=\pi/4$, and in Fig. 12 additional results for a completely incoherent state at $K=1$ and $\lambda=\pi/4$. We plot results from numerical simulations of the the KS model (1) and of our theory, Eq. (8) for the partially synchronized case and Eq. (10) for the incoherent state, respectively.
Figure 8: Covariance $R(\tau)$ and pseudo-covariance $\tilde{R}(\tau)$ of the complex order parameter fluctuation $\zeta(t)$ for a partially synchronized state at $K=5$ and $\lambda=\pi/4$ in the KS model (1). Numerical simulations (solid curve) vs. theoretical prediction (8) (dashed curve).
Figure 9: The same fluctuation characteristics as in Fig. 8 but for a partially synchronized state at $K=4$ and $\lambda=\pi/4$.
Figure 10: The same fluctuation characteristics as in Fig. 8 but for a partially synchronized state at $K=3$ and $\lambda=\pi/4$.
Figure 11: The same fluctuation characteristics as in Fig. 8 but for a partially synchronized state at $K=2$ and $\lambda=\pi/4$.
Figure 12: Covariance $R(\tau)$ and pseudo-covariance $\tilde{R}(\tau)$ of the complex order parameter fluctuation $\zeta(t)$ for a completely incoherent state at $K=1$ and $\lambda=\pi/4$ in the KS model (1). Numerical simulations (solid curve) vs. theoretical prediction (10) (dashed curve).
In addition, to demonstrate the effectiveness of our theoretical formulas for other system parameters, we present results for different values of the phase lag $\lambda$ and different frequency distributions $g(\omega)$. Fig. 13 shows how the variance $V$ of order parameter depends on the coupling strength $K$ for a unit variance Gaussian distribution (top row), for a triangular distribution (middle row)
$g(\omega)=\left\{\begin{array}[]{lcl}1-|\omega|&\mbox{ if }&|\omega|\leq 1,\\[2.84526pt] 0&\mbox{ if }&|\omega|>1,\end{array}\right$.
and for a Lorentzian distribution (bottom row)
$g(\omega)=\frac{1}{\pi(1+\omega^{2})},$
for several values of the phase lag $\lambda$. It is seen that the quality of our approximations does not significantly depend on neither the frequency distribution nor on the value of the phase lag.
Figure 13: Variance $V$ of the order parameter $r(t)$ as a function of the coupling strength $K$ for different values of the phase lag $\lambda$ for three frequency distributions $g(\omega)$: Gaussian (a)–(c), Triangular (d)–(f), and Lorentzian (g)–(i). Numerical simulations for the KS model (1) with $N=1,000$ (blue squares) vs. theoretical predictions (9) and (11) (solid curves).
Finally, in Figure 14 we show that the results for the full KS model (1) are indistinguishable by eye when a larger network with $10,000$ oscillators is simulated, indicative that $N=1,000$ are sufficient to study the thermodynamic limit.
Figure 14: Comparison of the results from numerical simulations of the KS model (1) for $N=1,000$ and $N=10,000$. (a)-(b): The solid curves show the covariance $R_{\mathrm{ps}}(\tau)$ of fluctuations $\zeta(t)$ of the modified order parameter $Z_{\mathrm{mod}}(t)$ for a partially synchronized state at $K=6$ and $\lambda=\pi/4$. Dashed curves show the theoretical prediction (8). (c): Variance of the order parameter $V$ as a function of the coupling strength $K$ (squares and crosses). The theoretical predictions (8) and (9) are shown as solid curves.
Appendix I Phase diffusion
Mean-field theory for partially synchronized states predicts that the argument of the complex order parameter
$Z(t)=\frac{1}{N}\sum\limits_{j=1}^{N}e^{i\psi_{j}(t)},$
uniformly increases or decreases with time. To subtract this drift, we can move into the corotating frame of reference and consider $\arg(Z(t)e^{-i\Omega t})$, where $\Omega$ is the average speed. However, this quantity experiences non-stationary irregular behaviour, resembling Brownian motion, as shown in Fig. 15. This phenomenon is known as phase diffusion [42, 40]. Due to this phase diffusion, the function $Z(t)e^{-i\Omega t}$ is difficult to characterize. Therefore, in our study we work with a modified order parameter
$Z_{\mathrm{mod}}(t)=Z(t)\frac{\overline{Z_{\mathrm{c}}(t)}}{|Z_{\mathrm{c}}(t)|},$
where
$Z_{\mathrm{c}}(t)=\frac{1}{N}\sum\limits_{j\in\mathcal{C}}e^{i\psi_{j}(t)}$
is the order parameter, defined as the sum over the coherent (locked) oscillators only. In numerical simulations, we identify the $j$th oscillator as coherent if its natural frequency $\omega_{j}$ satisfies the locking condition $|\omega_{j}-\Omega|\leq K\langle|Z(t)|\rangle$. It turns out, that the argument of the modified order parameter $Z_{\mathrm{mod}}(t)$ does not exhibit phase diffusion phenomenon, and the behavior of this complex function is nicely approximated by a two-dimensional Gaussian process [55]. This numerical observation is crucial for our analytical approach.
Figure 15: Phase diffusion in the KS model (1). The argument of the complex order parameter in the uniformly corotating frame (black curve) exhibits non-stationary random walk behavior. Contrary, the argument of the modified order parameter $Z_{\mathrm{mod}}(t)$ (red curve) exhibits stationary statistical behaviour. Shown are numerical results for the KS model (1) with $N=1,000$, $K=2$, $\lambda=\pi/4$, and a unit width Gaussian frequency distribution $g(\omega)$.
Appendix J Pairwise covariances of rogue oscillators
In Proposition 1, we derived an explicit formula for the solutions of the Adler equation (34). Simple calculations show that this formula has the equivalent form
$e^{i\theta(t)}e^{-i\theta(0)}=\frac{\displaystyle\chi_{+}e^{i\xi t}-\chi_{-}}{\displaystyle-\overline{\chi}_{-}e^{i\xi t}+\overline{\chi}_{+}}=\frac{\displaystyle\chi_{+}-\chi_{-}e^{-i\xi t}}{\displaystyle-\overline{\chi}_{-}+\overline{\chi}_{+}e^{-i\xi t}},$
where
$\chi_{-}=iHe^{-i\theta(0)}-(\Delta-\xi),\quad\chi_{+}=iHe^{-i\theta(0)}-(\Delta+\xi).$
We use this form to justify the following statements.
Proposition 2
If $\Delta>|H|$, we have $|\chi_{-}/\chi_{+}|<1$ and the following expansion holds
$e^{i\theta(t)}$ $=$ $-\frac{\displaystyle\chi_{-}}{\displaystyle\overline{\chi}_{+}}e^{i\theta(0)}+\frac{\displaystyle|\chi_{+}|^{2}-|\chi_{-}|^{2}}{\displaystyle\overline{\chi}_{+}\overline{\chi}_{-}}e^{i\theta(0)}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle\overline{\chi}_{-}}{\displaystyle\overline{\chi}_{+}}\right)^{n}e^{in\xi t}$ (38)
$=$ $\left\langle e^{i\theta(t)}\right\rangle-\frac{\displaystyle 2i\xi}{\displaystyle\overline{H}}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta-\xi}{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta+\xi}\right)^{n}e^{in\xi t},$
where $\left\langle e^{i\theta(t)}\right\rangle$ denotes the time average of $e^{i\theta(t)}$. On the other hand, if $\Delta<-|H|$, then $|\chi_{+}/\chi_{-}|<1$ and
$e^{i\theta(t)}$ $=$ $-\frac{\displaystyle\chi_{+}}{\displaystyle\overline{\chi}_{-}}e^{i\theta(0)}+\frac{\displaystyle|\chi_{-}|^{2}-|\chi_{+}|^{2}}{\displaystyle\overline{\chi}_{+}\overline{\chi}_{-}}e^{i\theta(0)}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle\overline{\chi}_{+}}{\displaystyle\overline{\chi}_{-}}\right)^{n}e^{-in\xi t}$
$=$ $\left\langle e^{i\theta(t)}\right\rangle+\frac{\displaystyle 2i\xi}{\displaystyle\overline{H}}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta+\xi}{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta-\xi}\right)^{n}e^{-in\xi t}.$
Proof: For brevity, we denote $\zeta=iHe^{-i\theta(0)}$. Then $e^{i\theta(0)}=i\overline{\zeta}/\overline{H}$. Next, using the definition $\xi=\sqrt{\Delta^{2}-|H|^{2}}$, we obtain
$|\chi_{+}|^{2}$ $=$ $(\zeta-\Delta-\xi)(\overline{\zeta}-\Delta-\xi)$
$=$ $(\Delta+\xi)(2\Delta-\zeta-\overline{\zeta}),$
$|\chi_{-}|^{2}$ $=$ $(\zeta-\Delta+\xi)(\overline{\zeta}-\Delta+\xi)$
$=$ $(\Delta-\xi)(2\Delta-\zeta-\overline{\zeta}),$
and
$\left(|\chi_{+}|^{2}-|\chi_{-}|^{2}\right)e^{i\theta(0)}$ $=$ $2\xi(2\Delta-\zeta-\overline{\zeta})i\overline{\zeta}/\overline{H}$
$=$ $-2i\xi(\overline{\zeta}^{2}-2\Delta\overline{\zeta}+|H|^{2})/\overline{H},$
$\overline{\chi}_{+}\overline{\chi}_{-}$ $=$ $(\overline{\zeta}-\Delta-\xi)(\overline{\zeta}-\Delta+\xi)$
$=$ $\overline{\zeta}^{2}-2\Delta\overline{\zeta}+|H|^{2}.$
Therefore,
$\frac{\displaystyle|\chi_{+}|^{2}-|\chi_{-}|^{2}}{\displaystyle\overline{\chi}_{+}\overline{\chi}_{-}}e^{i\theta(0)}=-\frac{\displaystyle 2i\xi}{\displaystyle\overline{H}}.$
On the other hand, we have
$\frac{\displaystyle\overline{\chi}_{-}}{\displaystyle\overline{\chi}_{+}}=\frac{\displaystyle\overline{\zeta}-\Delta+\xi}{\displaystyle\overline{\zeta}-\Delta-\xi}=\frac{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta-\xi}{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta+\xi}.$
Summarizing, this gives formula (38). The other formula in the proposition can be checked similarly.
Proposition 3
Suppose that $\theta_{j}(t)$ and $\theta_{k}(t)$ are two solutions of Eq. (34) with $\Delta=\Delta_{j}$ and $\Delta=\Delta_{k}$, respectively. The complex parameter $H$ in Eq. (34) is assumed to be the same for both indices $j$ and $k$. If $|\Delta_{j}|>|H|$ and $|\Delta_{k}|>|H|$, then
$\hskip-19.91692pt\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle-\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{-i\theta_{k}(t)}\rangle=$
$=\frac{\displaystyle 4\xi_{j}\>\mathrm{sgn}\>\Delta_{j}\>\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle|H|^{2}}\times$
$\times\sum\limits_{n,m=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}-\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}+\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}\right)^{n}\times$
$\times\left(\frac{\displaystyle iHe^{-i\theta_{k}(0)}-\Delta_{k}+\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle iHe^{-i\theta_{k}(0)}-\Delta_{k}-\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}\right)^{m}\times$
$\times e^{-im\xi_{k}\tau\>\mathrm{sgn}\>\Delta_{k}}\delta(n\xi_{j}\>\mathrm{sgn}\>\Delta_{j}-m\xi_{k}\>\mathrm{sgn}\>\Delta_{k}),$
and
$\hskip-19.91692pt\langle e^{i\theta_{j}(t)}e^{i\theta_{k}(t+\tau)}\rangle-\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{i\theta_{k}(t)}\rangle=$
$=-\frac{\displaystyle 4\xi_{j}\>\mathrm{sgn}\>\Delta_{j}\>\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle\overline{H}^{2}}\times$
$\times\sum\limits_{n,m=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}-\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}+\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}\right)^{n}\times$
$\times\left(\frac{\displaystyle i\overline{H}e^{i\theta_{k}(0)}+\Delta_{k}-\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle i\overline{H}e^{i\theta_{k}(0)}+\Delta_{k}+\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}\right)^{m}\times$
$\times e^{im\xi_{k}\tau\>\mathrm{sgn}\>\Delta_{k}}\delta(n\xi_{j}\>\mathrm{sgn}\>\Delta_{j}+m\xi_{k}\>\mathrm{sgn}\>\Delta_{k}),$
where $\xi_{j}=\sqrt{\Delta_{j}^{2}-|H|^{2}}$ and $\xi_{k}=\sqrt{\Delta_{k}^{2}-|H|^{2}}$.
Proof: The above formulas are straightforward consequences of Proposition 2 and the identity $\langle e^{i\nu t}\rangle=\delta(\nu)$
Remark 4
If $\xi_{j}$ and $\xi_{k}$ in Proposition 3 are incommensurable, then
$\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle=\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{-i\theta_{k}(t+\tau)}\rangle$
and
$\langle e^{i\theta_{j}(t)}e^{i\theta_{k}(t+\tau)}\rangle=\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{i\theta_{k}(t+\tau)}\rangle.$
Appendix K Daido’s theory
In [8], H. Daido proposed an analytical approach to approximate finite-size fluctuations in the KS model (1) with $\lambda=0$. In particular, for the subcritical case he obtained explicit expressions for the covariance function and the variance of the complex order parameter. We follow here his approach and determine expressions for $\lambda\neq 0$ in the subcritical case, and show that unlike for a Lorentz distribution of the natural frequencies this method is not well suited for a Gaussian frequency distribution.
Let us assume that the dynamics of the $i$th oscillator can be represented as
$\psi_{i}(t)=\Theta_{i}(t)+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\vartheta_{i}(t),$ (39)
where $\Theta_{i},\vartheta_{i}=\mathcal{O}(1)$ for $N\to\infty$. Upon substitution into the unnumbered equation below (1) in the main text, which we recall for convenience here,
$\dot{\psi}_{i}=\omega_{i}+K\>\mathrm{Im}\left(Z(t)e^{-i\psi_{i}}e^{-i\lambda}\right),$
we obtain
$\dot{\Theta}_{i}+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\dot{\vartheta}_{i}=\omega_{i}+K\>\mathrm{Im}\left(Z(t)e^{-i(\Theta_{i}+\vartheta_{i}/\sqrt{N})}e^{-i\lambda}\right).$
Next, assuming that $\langle Z(t)\rangle=0$ and hence $Z(t)=\mathcal{O}(1/\sqrt{N})$, and equating separately the terms of order $\mathcal{O}(1)$ and the terms of order $\mathcal{O}(1/\sqrt{N})$, we obtain
$\dot{\Theta}_{i}=\omega_{i}$ (40)
and
$\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\dot{\vartheta}_{i}=K\>\mathrm{Im}\left(Z(t)e^{-i\Theta_{i}}e^{-i\lambda}\right).$ (41)
Eq. (40) is solved by $\Theta_{i}(t)=\omega_{i}t+\Theta_{i}(0)$. Substituting this into Eq. (41), we obtain
$\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\dot{\vartheta}_{i}(t)=K\int_{0}^{t}\mathrm{Im}\left(Z(t^{\prime})e^{-i(\omega_{i}t^{\prime}+\Theta_{i}(0))}e^{-i\lambda}\right)dt^{\prime}$
with $\vartheta_{i}(0)=0$. Finally, substituting the expressions for the solution to Eqs. (40) and (41) into (39) and then substituting $\theta_{i}(t)=\psi_{i}(t)$ into the definition of the complex order parameter (2), we obtain a self-consistency equation for $Z(t)$:
$Z(t)$ $=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i(\omega_{i}t+\Theta_{i}(0))}\times$
$\times$ $\left(1+iK\int_{0}^{t}\mathrm{Im}\left(Z(t^{\prime})e^{-i(\omega_{i}t^{\prime}+\Theta_{i}(0))}e^{-i\lambda}\right)dt^{\prime}\right)$
$=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i(\omega_{j}t+\Theta_{j}(0))}$
$+$ $\frac{\displaystyle K}{\displaystyle 2}e^{-i\lambda}\int_{0}^{t}Z(t^{\prime})\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i\omega_{j}(t-t^{\prime})}dt^{\prime}$
- $\frac{\displaystyle K}{\displaystyle 2}e^{i\lambda}\int_{0}^{t}\overline{Z}(t^{\prime})\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i(\omega_{j}(t+t^{\prime})+2\Theta_{j}(0))}dt^{\prime},$ (43)
where we employed a Taylor expansion of the exponential function for $|Z(t)|\ll 1$.
In the next step, we will obtain an asymptotic formula for $Z(t)$ as $t\to\infty$. For this, we first notice that the function
$\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i(\omega_{j}(t+t^{\prime})+2\Theta_{j}(0))}$
is vanishing as $t\to\infty$ and hence can be neglected. Furthermore, for $N\gg 1$ we have
$\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i\omega_{j}t}\mapsto\int_{-\infty}^{\infty}g(\omega)e^{i\omega t}d\omega=\hat{g}(t).$
Thus, the complex order parameter $Z(t)$ given in (43) is the solution of the following Volterra integral equation
$Z(t)$ $=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i(\omega_{j}t+\Theta_{j}(0))}$ (44)
$+$ $\frac{\displaystyle K}{\displaystyle 2}e^{-i\lambda}\int_{0}^{t}\hat{g}(t-t^{\prime})Z(t^{\prime})dt^{\prime}.$
This equation can be solved using the Laplace transformation
$\mathcal{L}\>:\>f(t)\mapsto F(s)=\int_{0}^{\infty}e^{-st}f(t)dt.$
In particular, due to the convolution theorem, Eq. (44) we write
$(\mathcal{L}Z)(s)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\frac{\displaystyle e^{i\Theta_{j}(0)}}{\displaystyle s-i\omega_{j}}+\frac{\displaystyle K}{\displaystyle 2}e^{-i\lambda}(\mathcal{L}\hat{g})(s)(\mathcal{L}Z)(s),$
and therefore obtain
$(\mathcal{L}Z)(s)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\left(1-\frac{\displaystyle K}{\displaystyle 2}e^{-i\lambda}(\mathcal{L}\hat{g})(s)\right)^{-1}\frac{\displaystyle e^{i\Theta_{j}(0)}}{\displaystyle s-i\omega_{j}}.$
Now, assuming that
[…]
and that $1/Q(s)$ is a meromorphic function in the half-plane $\mathrm{Re}(s)<0$ that satisfies
$\lim\limits_{p\to\infty}\max\limits_{\pi/2\leq|\phi|\leq\pi}|1/Q(pe^{i\phi})|=0,$ (45)
we apply the inverse Laplace transformation and find that the long-term asymptotics of $Z(t)$ is given by
$Z(t)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\frac{\displaystyle e^{i(\omega_{j}t+\Theta_{j}(0))}}{\displaystyle Q(i\omega_{j})}.$
The covariance function of the complex order parameter $Z(t)$ is now evaluated as
$R_{\mathrm{Daido}}(\tau)$ $=$ $N\langle Z(t)\overline{Z(t+\tau)}\rangle$
$=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j,k=1}^{N}\frac{\displaystyle e^{-i(\omega_{k}\tau+\Theta_{k}(0)-\Theta_{j}(0))}}{\displaystyle Q(i\omega_{j})\overline{Q(i\omega_{k})}}\langle e^{i(\omega_{j}-\omega_{k})t}\rangle$
$=$ $\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}\frac{\displaystyle e^{-i\omega_{j}\tau}}{\displaystyle|Q(i\omega_{j})|^{2}}\approx\int_{-\infty}^{\infty}g(\omega)\frac{\displaystyle e^{-i\omega\tau}}{\displaystyle|Q(i\omega)|^{2}}d\omega,$
where we used $\langle e^{i(\omega_{j}-\omega_{k})t}\rangle=\delta_{jk}$. The variance of the order parameter $r(t)=|Z(t)|$ in the subcritical case is then given by
$V_{\mathrm{Daido}}=N\langle|Z(t)|^{2}\rangle-N\langle|Z(t)|\rangle^{2}=R_{\mathrm{Daido}}(0)-\frac{\displaystyle\pi}{\displaystyle 4},$ (46)
where we used our result (11) for $\langle|Z(t)|\rangle$.
Figure 16: Variance $V$ of the order parameter $r(t)$ versus coupling strength $K$ in the subcritical case of the KS model (1) with $\lambda=\pi/4$ and a Gaussian distribution. Theoretical predictions from our approach (11) (thick line) and from Daido’s theory (46) (circles) are compared with the results of numerical simulations with $N=1,000$ phase oscillators (squares). The vertical dashed line indicates the critical coupling strength.
Although the above formulas have been shown to be in good agreement with the numerical observations obtained from the full KS model (1) with a Lorentzian frequency distribution and $\lambda=0$, in the case of a Gaussian frequency distribution they fail to reproduce the observed behavior as seen in Fig. 16. In particular, they can be worse than our $(K,\lambda)$-independent formulas (10) and (11). Note that condition (45) is not satisfied for a Gaussian frequency distribution.